# AFA-LoRA: Enabling Non-Linear Adaptations in LoRA with Activation Function Annealing

### v1

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
The growth of Large Language Models (LLMs)(Vaswani et al., 2017; Team et al., 2025; Achiam et al., 2023; Liu et al., 2024a; Bai et al., 2023) has revolutionized natural language processing. However, due to the extremely large number of parameters in these massive models, fully fine-tuning them for downstream tasks is usually infeasible. This challenge led to the development of the Parameter Efficient Fine-Tuning (PEFT) method, which aims to achieve competitive performance by training only a small fraction of the model’s parameters. Among these, Low-Rank Adaptation (LoRA) (Hu et al., 2022) has become a major method. LoRA freezes the weights $W$ of the pre-trained model and inserts a trainable rank decomposition matrix (adapter) into each layer. The update is parameterized as $\Delta W=\frac{\alpha}{r}BA$, where $A\in\mathbb{R}^{r\times d_{\text{in}}}$ and $B\in\mathbb{R}^{d_{\text{out}}\times r}$ are the low-rank matrices with rank $r$, and $\alpha$ is a scaling hyperparameter. This design allows the adapter to be seamlessly merged back into the main model after training. ($W_{\text{new}}=W+\frac{\alpha}{r}BA$).
@@
We evaluated AFA-LoRA on a variety of tasks. In the Supervised Fine-Tuning (SFT) benchmark, it narrowed the performance gap between standard LoRA and fully parametric fine-tuning. Secondly, we integrated AFA-LoRA into the GRPO framework(Shao et al., 2024) for reinforcement learning, demonstrating that it also effectively reduces the gap between GRPO-LoRA and full-parameter GRPO, showing that it works well for more than just SFT. We also integrated AFA-LoRA into the draft model in Eagle, a popular speculative decoding framework. We added LoRA adapters to the draft model in Eagle and trained them together with the main weights. This enabled the draft model to accept longer token sequences, showing that AFA-LoRA can be well adapted to different tasks.
@@ 3.1 Parameter-Efficient Fine-Tuning
The high cost of full fine-tuning has led to the creation of PEFT-based approaches. Early approaches include adapter-based methods, which insert small, trainable modules between layers of a pre-trained model (Houlsby et al., 2019), and prompt-based techniques like prompt-tuning and prefix-tuning, which optimize continuous input vectors (Lester et al., 2021; Li & Liang, 2021). Among these methods, Low-Rank Adaptation (LoRA) (Hu et al., 2022) is widely used because it is efficient, and its adapters can be easily merged into the main model after training; thus, there is no extra cost during inference. In our method, we use LoRA as a starting point and focus on improving its ability to learn complex patterns while retaining the integrable features.
@@ 3.2 Advances in the LoRA Framework
The success of LoRA has inspired extensive research aimed at improving its efficiency and performance. One approach is adaptive parameter allocation (e.g. AdaLoRA) (Zhang et al., 2023). AdaLoRA changes the rank of LoRA matrices during training to better utilize computational resources. Other innovative methods include VeRA (Kopiczko et al., 2023), which focuses on parameter reduction, and DoRA (Liu et al., 2024b), which decouples the magnitude and direction of weight updates. QLoRA (Dettmers et al., 2023) allows for fine-tuning large models using quantized weights, which saves memory and speeds up training.
@@ 3.3 Activation Functions in Model Adaptation
Previous work such as PReLU (He et al., 2015) has shown that allowing the activation function to change during training is feasible. Our method follows this idea by gradually changing the non-linearity in adapters over time.
@@ 4.2.1 Annealed Activation Function
Where, $\beta$ is a value that decreases from 1 to 0 as training continues. This means that when $\beta=1$, we will adapt de non-linear function $\sigma(x)$ for updating, and when $\beta=0$, we simply use $x$.
@@ 4.2.2 AFA-LoRA Forward Computation
$\begin{split}h&=W_{0}x+B\cdot\phi(Ax;\beta(t))\\ &=W_{0}x+B\cdot\left[\beta(t)\cdot\sigma(Ax)+(1-\beta(t))\cdot Ax\right].\end{split}$ (4)
@@ 4.4 Training Algorithm
Input: Model $W_{0}$, Dataset $D$, Total steps $T$, Annealing range $T_{\text{start}},T_{\text{end}}]$
 Initialize LoRA parameters $\theta=\{A,B\}$.
for step $t=1$ in $T$ do
  $\beta(t)\leftarrow\max(0,1-(t-T_{\text{start}})/(T_{\text{end}}-T_{\text{start}}))$
  Sample batch $(x,y)\sim D$
  Forward pass: $h=W_{0}x+B\cdot[\beta(t)\sigma(Ax)+(1-\beta(t))Ax]$
  Compute loss $\mathcal{L}$
  Update $\theta$ via gradient descent
end for
 Merge adapter: $W^{\prime}\leftarrow W_{0}+BA$
4.5 Application to Reinforcement Learning
GRPO (Shao et al., 2024) is a widely used framework for efficient policy optimization in reinforcement learning. The GRPO-LoRA framework uses LoRA to make reinforcement learning fine-tuning more efficient. Built upon this conclusion, we introduce AFA-LoRA layers to the policy network to present a potential application of AFA-LoRA. This means that during training, the value $\beta(t)$ changes over time and affects how each layer adapts. As $\beta$ approaches zero, the policy becomes more stable and focuses on exploitation. Importantly, even though behavior changes during training, the main GRPO objective remains the same. After training finishes, we can merge all adapters back into the main model, just like with standard LoRA, no extra cost happens during inference time. We present experimental results in the following content.
Table 1: Commonsense reasoning evaluation results (Accuracy %) on Llama-3-8B. LoRA and DORA are shown in separate groups for clear comparison. Results show the 30% decay configuration for each AFA placement variant. Avg is the macro-average across all 8 tasks. The best result for each method is bold.
FULL-SFT 82.94 92.68 74.98 96.71 89.40 90.26 83.21 86.11 87.07
LoRA Baseline
LORA 80.03 90.99 74.56 96.03 87.00 88.63 81.17 86.11 85.57
AFA-LoRA A-$\sigma$-B 79.78 91.25 75.60 96.03 86.80 88.52 81.37 86.66 85.75
AFA-LoRA A-B-$\sigma$ 81.66 91.46 74.80 95.66 87.20 89.06 81.99 86.42 86.03
AFA-LoRA $\sigma$-A-$\sigma$-B 80.38 90.91 74.95 96.11 87.40 88.52 80.04 87.45 85.72
AFA-LoRA A-$\sigma$-B-$\sigma$ 80.72 91.41 75.08 95.87 86.40 89.17 81.06 87.37 85.89
AFA-LoRA $\sigma$-A-B-$\sigma$ 81.14 91.33 75.44 95.71 88.40 88.68 81.53 87.06 86.16
AFA-LoRA $\sigma$-A-$\sigma$-B-$\sigma$ 80.29 90.82 75.96 95.74 87.20 89.06 80.60 85.87 85.69
DORA Baseline
DORA 80.46 90.45 75.66 95.77 85.00 87.92 81.06 87.29 85.45
AFA-DORA A-$\sigma$-B 79.69 91.25 75.93 95.80 86.40 88.74 80.45 86.74 85.63
AFA-DORA A-B-$\sigma$ 80.97 91.62 75.56 95.90 88.40 89.11 82.59 86.58 86.34
AFA-DORA $\sigma$-A-$\sigma$-B 80.20 91.07 75.99 96.05 87.60 89.33 81.16 87.21 86.08
AFA-DORA A-$\sigma$-B-$\sigma$ 79.52 90.74 75.35 95.75 85.40 89.88 81.21 87.05 85.61
AFA-DORA $\sigma$-A-B-$\sigma$ 81.48 91.58 76.17 95.66 86.80 89.00 81.42 86.97 86.14
AFA-DORA $\sigma$-A-$\sigma$-B-$\sigma$ 81.65 91.07 76.60 95.70 87.00 88.57 81.26 87.21 86.13
4.6 Application to Speculative Decoding
To prove the flexibility of our method, we also adapt AFA-LoRA to speculative decoding frameworks like EAGLE (Li et al., 2024). For each linear layer in the draft model, we add a LoRA adapter and train both the main weights and LoRA parameters together using activation function annealing. The forward calculation for each layer is:
$y=W_{\mathrm{main}}\,x+B\,[\,\beta\,\sigma(Ax)\,+\,(1{-}\beta)\,Ax]$ (6)
where $W_{\mathrm{main}}$ denotes pre-trained weights; $(A,B)$ are low-rank matrices forming the LoRA branch; $\sigma(\cdot)$ is a non-linear activation function; and $\beta$ linearly decays from 1 to 0 throughout training. At inference time (${\beta}=0$), all adapters are merged into base weights via simple addition without any extra latency. As shown in Section 5.3, this method increases token acceptance rates and maintains the speed benefits of speculative decoding.
@@ 5.1 Supervised Fine-Tuning Experiments
We include both LoRA and DoRA (Weight-Decomposed Low-Rank Adaptation) (Liu et al., 2024b) as baselines in our experiments. DoRA is an advanced parameter-efficient fine-tuning method that splits weight updates into direction and magnitude, enabling it to capture more complex changes than standard LoRA.
@@
We implemented activation function annealing with LoRA and DoRA on supervised fine-tuning with the Llama-3-8B (AI@Meta, 2024) model and the Commonsense-170K (Hu et al., 2023) dataset, which is designed for commonsense reasoning. We evaluate performance across eight benchmarks: ARC-Challenge (Clark et al., 2018), ARC-Easy, BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OpenBookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), Social IQA (Sap et al., 2019), and WinoGrande (Sakaguchi et al., 2021). Our experiments compare AFA-enhanced methods with three baselines: full parameter fine-tuning (Full-SFT), standard LoRA, and DoRA. For both LoRA and DoRA, we try seven different ways of placing the annealed activation function in the adapter structure, using rank $r=32$ and scaling factor $\alpha=64$. The value of $\beta$ decreases linearly over the first 30% of training steps, so in Equation (5), we set $T_{start}=0$ and $T_{end}=\frac{T}{3}$.
5.2 Reinforcement Learning Experiments
We conduct reinforcement learning experiments using the GRPO (Group Relative Policy Optimization) framework, implemented with the Verl training system. This method uses preference optimization within groups of responses to keep policy updates stable. We test Qwen2.5-Instruct models of different sizes (3B, 7B, 14B, and 32B)(Team, 2024; Yang et al., 2024) on GSM8K(Cobbe et al., 2021), a dataset for math reasoning tasks. For training, we use the AdamW optimizer with model-specific learning rates ($2\times 10^{-5}$ for 3B; $1.5\times 10^{-5}$ for 7B/14B; $1.3\times 10^{-5}$ for 32B). The global batch size is set to 1024 and is split into PPO mini-batches of 256 samples for each global batch. Each sequence has up to 512 prompt tokens and up to 1024 response tokens. The loss includes KL divergence regularization ($\beta=0.001$), but no entropy term is used. For each prompt, five responses are generated using vLLM (Kwon et al., 2023). We adapt gradient checkpointing and FSDP to save memory and speed up training (Zhao et al., 2023). All experiments are carried out with a total of fifteen epochs. All variants of LoRA are set with rank $r=64$ and scaling factor $\alpha=32$. For AFA-LoRA, $\beta$ decays linearly over the first $30\%$ steps in all tested placements.
@@ 5.3 Speculative Sampling Experiments
Here, $A$ and $B$ are small trainable matrices, and $\sigma(\cdot)$ are activation functions, such as ReLU or GeLU. The value of $\beta$ is initialized as 1 (completely non-linear) at the beginning of training. As the training continues, $\beta$ will decay to 0, which means that the transformation converge to linear form.
By applying AFA-LoRA adapters and training them together with the main weights, the draft model better fit the training data significantly. This works considerably well because the draft model is much smaller than large language models, this extra flexibility makes a significant difference in performance. Our experiments show that using AFA-LoRA in this way allows the model to learn faster and achieve better results.
After training finishes, all these extra parameters can be merged into the weights of original draft model easily. This implies that no side effect will exist during the inference phase for generatintg text.
@@
As shown in Table 3, integrating activation function annealing improves speculative decoding performance across all evaluation benchmarks. The SiLU(Elfwing et al., 2018) annealing strategy achieves the highest acceptance rates in most scenarios, with particularly notable gains on the HumanEval benchmark, where it improves acceptance by 1.0% in chain mode and 0.8% in tree mode compared to the pure LoRA baseline.
@@
Overall, adding AFA-LoRA adapters contributes to performance gain of Eagle in accepting tokens during speculative decoding. This shows that our method is a strong choice for improvin draft models in real-world generation tasks.
@@ 6 Ablation Studies
In this section, we explore the impact of different annealing schedules and activation functions on AFA-LoRA performance. While our main experiments (Section 5) focused on ReLU activation with a 30% decay duration, here we evaluate SiLU and GeLU activations across three decay schedules. We performed ablation experiments using the same Llama3-8B LoRA SFT task described in Section  5.1, keeping the same hyperparameters (rank $r=32$, scaling factor $\alpha=64$) and evaluation benchmarks. The key difference lies in the annealing configurations: instead of changing the position of activation function as in the table  1, we fix the position with the manner of A-$\sigma$-B and change the activation function type ($\sigma\in\{\text{SiLU},\text{GeLU}\}$) and the decay schedules ($T_{\text{end}}\in\{0.3T,0.6T,T\}$).

### v2

@@ Abstract
rmTeXGyreTermesX [*devanagari]rmLohit Devanagari [*arabic]rmNoto Sans Arabic
AFA-LoRA: Enabling Non-Linear Adaptations in LoRA with Activation Function Annealing
Jiacheng Li1thanks: These authors contributed equally to this work., Jianchao Tan [^1] footnotemark: 1, Zhidong Yang2, Feiye Huo1, Yerui Sun1, Yuchen Xie1, Xunliang Cai1, 1Meituan, Beijing, China 2Hong Kong University of Science and Technology, Hong Kong SAR, China lijiacheng14@meituan.com
@@ 1 Introduction
The growth of Large Language Models (LLMs)Vaswani et al. (2017); Team et al. (2025); Achiam et al. (2023); Liu et al. (2024a); Bai et al. (2023) has revolutionized natural language processing. However, due to the extremely large number of parameters in these massive models, fully fine-tuning them for downstream tasks is usually infeasible. This challenge led to the development of the Parameter Efficient Fine-Tuning (PEFT) method, which aims to achieve competitive performance by training only a small fraction of the model’s parameters. Among these, Low-Rank Adaptation (LoRA) Hu et al. (2022) has become a major method. LoRA freezes the weights $W$ of the pre-trained model and inserts a trainable rank decomposition matrix (adapter) into each layer. The update is parameterized as $\Delta W=\frac{\alpha}{r}BA$, where $A\in\mathbb{R}^{r\times d_{\text{in}}}$ and $B\in\mathbb{R}^{d_{\text{out}}\times r}$ are the low-rank matrices with rank $r$, and $\alpha$ is a scaling hyperparameter. This design allows the adapter to be seamlessly merged back into the main model after training. ($W_{\text{new}}=W+\frac{\alpha}{r}BA$).
@@
We evaluated AFA-LoRA on a variety of tasks. In the Supervised Fine-Tuning (SFT) benchmark, it narrowed the performance gap between standard LoRA and fully parametric fine-tuning. Secondly, we integrated AFA-LoRA into the GRPO frameworkShao et al. (2024) for reinforcement learning, demonstrating that it also effectively reduces the gap between GRPO-LoRA and full-parameter GRPO, showing that it works well for more than just SFT. We also integrated AFA-LoRA into the draft model in Eagle, a popular speculative decoding framework. We added LoRA adapters to the draft model in Eagle and trained them together with the main weights. This enabled the draft model to accept longer token sequences, showing that AFA-LoRA can be well adapted to different tasks.
@@ 3.1 Parameter-Efficient Fine-Tuning
The high cost of full fine-tuning has led to the creation of PEFT-based approaches. Early approaches include adapter-based methods, which insert small, trainable modules between layers of a pre-trained model Houlsby et al. (2019), and prompt-based techniques like prompt-tuning and prefix-tuning, which optimize continuous input vectors Lester et al. (2021); Li and Liang (2021). Among these methods, Low-Rank Adaptation (LoRA) Hu et al. (2022) is widely used because it is efficient, and its adapters can be easily merged into the main model after training; thus, there is no extra cost during inference. In our method, we use LoRA as a starting point and focus on improving its ability to learn complex patterns while retaining the integrable features.
@@ 3.2 Advances in the LoRA Framework
The success of LoRA has inspired extensive research aimed at improving its efficiency and performance. One approach is adaptive parameter allocation (e.g. AdaLoRA) Zhang et al. (2023). AdaLoRA changes the rank of LoRA matrices during training to better utilize computational resources. Other innovative methods include VeRA Kopiczko et al. (2023), which focuses on parameter reduction, and DoRA Liu et al. (2024b), which decouples the magnitude and direction of weight updates. QLoRA Dettmers et al. (2023) allows for fine-tuning large models using quantized weights, which saves memory and speeds up training.
@@ 3.3 Activation Functions in Model Adaptation
Previous work such as PReLU He et al. (2015) has shown that allowing the activation function to change during training is feasible. Our method follows this idea by gradually changing the non-linearity in adapters over time.
@@ 4.2.1 Annealed Activation Function
Where, $\beta$ is a value that decreases from 1 to 0 as training continues. This means that when $\beta=1$, we will adapted non-linear function $\sigma(x)$ for updating, and when $\beta=0$, we simply use $x$.
@@ 4.2.2 AFA-LoRA Forward Computation
$\begin{split}h&=W_{0}x+B\phi(Ax;\beta(t))\\ &=W_{0}x+B\left[\beta(t)\sigma(Ax)+(1-\beta(t))Ax\right]\end{split}$ (4)
@@ 4.4 Training Algorithm
1:Input: Model $W_{0}$, Dataset $D$, Total steps $T$, Annealing range $T_{\text{start}},T_{\text{end}}]$
2:Initialize LoRA parameters $\theta=(A,B)$.
3:for $t=1$ to $T$ do
4:  $\beta(t)=1-(t-T_{\mathrm{{start}}})/(T_{\mathrm{{end}}}-T_{\mathrm{{start}}})$
5:  Sample batch $(x,y)\penalty 10000\ {\sim}\penalty 10000\ D$
6:  Forward pass: $h=W_{0}x+B[\beta(t)\sigma(Ax)+(1-\beta(t))Ax]$
7:  Compute loss $\mathcal{L}$
8:  Update $\theta$ via gradient descent
9:end for
10:Merge adapter: $W^{\prime}=W_{0}+BA$
@@ 5.1 Supervised Fine-Tuning Experiments
Table 1: Commonsense reasoning evaluation results (Accuracy %) on Llama-3-8B. Results show the 30% decay configuration for each AFA placement variant. Avg is the macro-average across all 8 tasks. The best result for each method is bold.
FULL-SFT 82.94 92.68 74.98 96.71 89.40 90.26 83.21 86.11 87.07
LoRA Baseline 80.03 90.99 74.56 96.03 87.00 88.63 81.17 86.11 85.57
A-$\sigma$-B 79.78 91.25 75.60 96.03 86.80 88.52 81.37 86.66 85.75
A-B-$\sigma$ 81.66 91.46 74.80 95.66 87.20 89.06 81.99 86.42 86.03
$\sigma$-A-$\sigma$-B 80.38 90.91 74.95 96.11 87.40 88.52 80.04 87.45 85.72
A-$\sigma$-B-$\sigma$ 80.72 91.41 75.08 95.87 86.40 89.17 81.06 87.37 85.89
$\sigma$-A-B-$\sigma$ 81.14 91.33 75.44 95.71 88.40 88.68 81.53 87.06 86.16
$\sigma$-A-$\sigma$-B-$\sigma$ 80.29 90.82 75.96 95.74 87.20 89.06 80.60 85.87 85.69
DORA Baseline 80.46 90.45 75.66 95.77 85.00 87.92 81.06 87.29 85.45
A-$\sigma$-B 79.69 91.25 75.93 95.80 86.40 88.74 80.45 86.74 85.63
A-B-$\sigma$ 80.97 91.62 75.56 95.90 88.40 89.11 82.59 86.58 86.34
$\sigma$-A-$\sigma$-B 80.20 91.07 75.99 96.05 87.60 89.33 81.16 87.21 86.08
A-$\sigma$-B-$\sigma$ 79.52 90.74 75.35 95.75 85.40 89.88 81.21 87.05 85.61
$\sigma$-A-B-$\sigma$ 81.48 91.58 76.17 95.66 86.80 89.00 81.42 86.97 86.14
$\sigma$-A-$\sigma$-B-$\sigma$ 81.65 91.07 76.60 95.70 87.00 88.57 81.26 87.21 86.13
We include both LoRA and DoRA (Weight-Decomposed Low-Rank Adaptation) Liu et al. (2024b) as baselines in our experiments. DoRA is an advanced parameter-efficient fine-tuning method that splits weight updates into direction and magnitude, enabling it to capture more complex changes than standard LoRA.
We implemented activation function annealing with LoRA and DoRA on supervised fine-tuning with the Llama-3-8B AI@Meta (2024) model and the Commonsense-170K Hu et al. (2023) dataset, which is designed for commonsense reasoning. We evaluate performance across eight benchmarks: ARC-Challenge Clark et al. (2018), ARC-Easy, BoolQ Clark et al. (2019), HellaSwag Zellers et al. (2019), OpenBookQA Mihaylov et al. (2018), PIQA Bisk et al. (2020), Social IQA Sap et al. (2019), and WinoGrande Sakaguchi et al. (2021). Our experiments compare AFA-enhanced methods with three baselines: full parameter fine-tuning (Full-SFT), standard LoRA, and DoRA. For both LoRA and DoRA, we try seven different ways of placing the annealed activation function in the adapter structure, using rank $r=32$ and scaling factor $\alpha=64$. The value of $\beta$ decreases linearly over the first 30% of training steps, so in Equation (5), we set $T_{start}=0$ and $T_{end}=\frac{T}{3}$.
5.2 Reinforcement Learning Experiments
@@
We conduct reinforcement learning experiments using the GRPO (Group Relative Policy Optimization) framework, implemented with the Verl training system. This method uses preference optimization within groups of responses to keep policy updates stable. We test Qwen2.5-Instruct models of different sizes (3B, 7B, 14B, and 32B)Team (2024); Yang et al. (2024) on GSM8KCobbe et al. (2021), a dataset for math reasoning tasks. For training, we use the AdamW optimizer with model-specific learning rates ($2\times 10^{-5}$ for 3B; $1.5\times 10^{-5}$ for 7B/14B; $1.3\times 10^{-5}$ for 32B). The global batch size is set to 1024 and is split into PPO mini-batches of 256 samples for each global batch. Each sequence has up to 512 prompt tokens and up to 1024 response tokens. The loss includes KL divergence regularization ($\beta=0.001$), but no entropy term is used. For each prompt, five responses are generated using vLLM Kwon et al. (2023). We adapt gradient checkpointing and FSDP to save memory and speed up training Zhao et al. (2023). All experiments are carried out with a total of fifteen epochs. All variants of LoRA are set with rank $r=64$ and scaling factor $\alpha=32$. For AFA-LoRA, $\beta$ decays linearly over the first $30\%$ steps in all tested placements.
@@ 5.3 Speculative Sampling Experiments
where $A,B$ are trainable low-rank matrices, $\sigma$ is an activation function (ReLU/GeLU/SiLU), and $\beta$ decays from 1 to 0 during training.
Joint training with AFA-LoRA adapters enhances the draft model’s capacity, particularly beneficial for smaller models like Eagle. Post-training, adapters merge into base weights without inference overhead.
@@
SiLU Elfwing et al. (2018) annealing yields the highest acceptance rates (Table 3), with +1.0%/+0.8% gains on HumanEval for chain/tree decoding versus LoRA baseline.
@@
Overall, adding AFA-LoRA adapters contributes to performance gain of Eagle in accepting tokens during speculative decoding. This shows that our method is a strong choice for improving draft models in real-world generation tasks.
@@ 6 Ablation Studies
In this section, we explore the impact of different annealing schedules and activation functions on AFA-LoRA performance. While our main experiments (Section 5) focused on ReLU activation with a 30% decay duration, here we evaluate SiLU and GeLU activations across three decay schedules. We performed ablation experiments using the same Llama3-8B LoRA SFT task described in Section 5.1, keeping the same hyperparameters (rank $r=32$, scaling factor $\alpha=64$) and evaluation benchmarks. The key difference lies in the annealing configurations: instead of changing the position of activation function as in the table 1, we fix the position with the manner of A-$\sigma$-B and change the activation function type ($\sigma\in\{\text{SiLU},\text{GeLU}\}$) and the decay schedules ($T_{\text{end}}\in\{0.3T,0.6T,T\}$).
@@ 7 Conclusion
8 Limitations
While AFA-LoRA performed well across multiple tasks, this study still has some limitations that warrant further investigation in future work.
AFA-LoRA introduces additional hyperparameters, including annealing time ($T_{\text{start}}$, $T_{\text{end}}$), activation function type (ReLU, SiLU, GeLU), and their placement in the adapter. While experiments show that 30% annealing time and SiLU activation perform well on most tasks, these optimal configurations may vary across different tasks, datasets, or model sizes, requiring task-specific tuning. Additionally, the continuous gradient variations during annealing may affect training stability, particularly in large-scale distributed training or resource-limited environments. Future work could explore adaptive annealing strategies, more stable optimization methods, and systematic evaluations under various hardware setups to enhance the method’s robustness and practicality.